40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17165     Entry time: Thu Sep 29 18:01:14 2022     In reply to: 17157     Reply to this: 17166
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BH55 LSC Model Updates - part IV 

More model changes

c1lsc:

  • BH55_I and BH55_Q are now being read at ADC_0_14 and ADC_0_15. The ADC_0_20 and ADC_0_21 are bad due to faulty whitening filter board.
  • The whitening switch controls were also shifted accordingly.
  • the slow epics channels for BH55 anti-aliasing switch and whitening switch were added in /cvs/cds/caltech/target/c1iscaux/C1_ISC-AUX_LSCPDs.db

c1mcs:

  • MC1, MC2, and MC3 are running on new suspension models now.

c1hpc:

  • DCPD_A and DCPD_B have been renamed to BHDC_A and BHDC_B following naming convention at other ports.
  • After the input summing matrix, the signals are called BHDC_SUM and BHDC_DIFF now.
  • BHDC_SUM and BHDC_DIFF can be directly using in sensing matrix bypassing the dither demodulation (to be used for DC locking)
  • BH55_I and BH55_Q are also sent for dither demodulation now (to be used in double dither method, RF and audio).
  • SHMEM channel names to c1bac were changed.

c1bac:

  • Conformed with new SHMEM channel names from c1hpc
ELOG V3.1.3-